# Improving Diffusion Inverse Problem Solving with Decoupled Noise Annealing

### v1

@@ 3.1 Posterior Sampling with Decoupled Noise Annealing
$p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t},{\mathbf{y}})=\frac{p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})p({\mathbf{y}}\mid{\mathbf{x}}_{0},{\mathbf{x}}_{t})}{p({\mathbf{y}}\mid{\mathbf{x}}_{t})}\propto p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})p({\mathbf{y}}\mid{\mathbf{x}}_{0}).$ (6)
To sample ${\mathbf{x}}_{0\mid{\mathbf{y}}}$ from this unnormalized distribution, we propose to run Langevin dynamics[38], given by
${\mathbf{x}}_{0}^{(j+1)}={\mathbf{x}}_{0}^{(j)}+\eta\cdot\left(\nabla_{{\mathbf{x}}_{0}^{(j)}}\log p({\mathbf{x}}_{0}^{(j)}\mid{\mathbf{x}}_{t})+\nabla_{{\mathbf{x}}_{0}^{(j)}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{0}^{(j)})\right)+\sqrt{2\eta}\bm{\epsilon}_{j},$ (7)
Following previous works [3, 22, 23], we approximate the conditional distribution $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})$ with a Gaussian distribution
where $\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})$ is an estimator of ${\mathbf{x}}_{0}$ given ${\mathbf{x}}_{t}$, and the variance $r_{t}^{2}$ is specified using heuristics. Given a pre-trained diffusion model ${\bm{s}}_{\bm{\theta}}({\mathbf{x}}_{t},t)$, one way to compute $\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})$ is to solve the (unconditional) probability flow ODE starting at ${\mathbf{x}}_{t}$. We leave details in Section D.2. When the measurement noise is an isotropic Gaussian, i.e., ${\mathbf{n}}\sim\mathcal{N}({\bm{0}},\beta_{\mathbf{y}}^{2}{\bm{I}})$, the update rule simplifies to:
@@
${\mathbf{x}}_{0}^{(j+1)}={\mathbf{x}}_{0}^{(j)}-\eta\cdot\nabla_{{\mathbf{x}}_{0}^{(j)}}\left(\frac{\|{\mathbf{x}}_{0}^{(j)}-\hat{{\mathbf{x}}}_{0}({\mathbf{x}}_{t})\|^{2}}{2r_{t}^{2}}+\frac{\|{\mathcal{A}}({\mathbf{x}}_{0}^{(j)})-{\mathbf{y}}\|^{2}}{2\beta_{\mathbf{y}}^{2}}\right)+\sqrt{2\eta}\bm{\epsilon}_{j},$ (9)
Combining Eq. 9 with Proposition 1, we define an algorithm called Decoupled Annealing Posterior Sampling (DAPS), summarized in Algorithm 1. In particular, given a noise schedule $\sigma_{t}$ and a time discretization $\{t_{i},i=0,\dots,N_{A}\}$, we iteratively sample ${\mathbf{x}}_{t_{i}}$ from the measurement-conditioned time-marginal $p({\mathbf{x}}_{t_{i}}\mid{\mathbf{y}})$ for $i=N_{A},N_{A}-1,\dots,0$, following Eq. 5 in Proposition 1. This algorithm provides an approximate sample ${\mathbf{x}}_{0}$ from the posterior distribution $p({\mathbf{x}}_{0}\mid{\mathbf{y}})$.
The computational cost of Langevin dynamics mostly originates from evaluating the measurement function $\mathcal{A}$. In most image restoration tasks, this operation is significantly more efficient compared to evaluating the diffusion model. As demonstrated in Section C.1, Langevin dynamics introduce only a small overhead to the sampling process. Moreover, this framework can be adapted to sampling with pre-trained latent diffusion models by factorizing the probabilistic graphical model for the latent diffusion process, as shown in Fig. 3(b). Further details are provided in Appendix A.
3.2 Discussion and Connection with Existing Methods
Comparison with other posterior sampling methods. Unlike most previous algorithms, our sampling method does not solve a specific SDE/ODE; instead, we recursively sample from the time-marginal $p({\mathbf{x}}_{t}\mid{\mathbf{y}})$ with noise annealing to zero. We therefore decouple the dependency of ${\mathbf{x}}_{t}$ and ${\mathbf{x}}_{t+\Delta t}$ in the sampling process. We argue that decoupling helps correct the errors accumulated in early diffusion sampling steps by allowing non-local transitions. This is particularly important when the measurement function is nonlinear. As a concrete example, Fig. 4 compares DAPS and DPS when solving a 2D nonlinear inverse problem with a Gaussian mixture prior. We visualize the trajectories of ${\mathbf{x}}_{t}$ from time $T$ to 00 for both methods. For DAPS, points on the trajectory have significantly larger variations compared to those of DPS. As a result, DPS converges to wrong solutions, but DAPS is able to approximate the true posterior distribution. We provide more discussions in Appendix E.
Comparison with optimization-based methods. Although our method does not directly involve inner-loop optimization, we highlight that it has connections with existing optimization-based solvers for inverse problems. For example, ReSample [25] alternate between denoising, optimizing, and resampling to solve inverse problems using latent diffusion models. Our method also resembles  [25] if we set standard deviation $\beta_{\mathbf{y}}\to 0$ and assume $\eta\beta_{\mathbf{y}}^{2}$ is constant in Eq. 12.
4 Experiments
4.1 Experimental Setup
As mentioned in Section 3, we implement a few-step Euler ODE solver to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, while maintaining the same number of neural function evaluations (NFE) for different noise levels. In our experiments, we use DAPS-1K for all linear tasks and DAPS-4K for all nonlinear tasks. DAPS-1K uses $4$ ODE solver NFE and $250$ annealing scheduling steps, while DAPS-4K uses $10$ ODE solver NFE and $400$ annealing scheduling steps. Ablation studies on their effects are shown in Section 4.3. The corresponding latent version, LatentDAPS, follows the same settings but performs sampling in the latent space of VAEs. We use $100$ Langevin steps per denoising iteration and tune learning rates separately for each task. Further details on model configurations, samplers, and other hyperparameters are provided in Appendix D, along with more discussion on sampling efficiency in Section C.1.
Datasets and metrics. Adopting the previous convention, we test our method on two image datasets, FFHQ $256\times 256$ [41] and ImageNet $256\times 256$ [42]. To evaluate our method, we use $100$ images from the validation set for both FFHQ and ImageNet. We include Learned Perceptual Image Patch Similarity (LPIPS) score [43] and peak signal-to-noise-ratio (PSNR) as our main evaluation metrics. For both our method and baselines, we use the versions implemented in piq [44] with all images normalized to the range $[0,1]$. The replace-pooling option is enabled for LPIPS evaluation.
We consider three nonlinear inverse problems: (1) phase retrieval, (2) high dynamic range (HDR) reconstruction, and (3) nonlinear deblurring. Due to the inherent instability of phase retrieval, we adopt the strategy from DPS [3], using an oversampling rate of 2.0 and reporting the best result out of four independent samples. The goal of HDR reconstruction is to recover a higher dynamic range image (factor of 2) from a low dynamic range image. For nonlinear deblurring, we use the default setting as described in [45]. All linear and nonlinear measurements are subject to white Gaussian noise with a standard deviation of $\beta_{\mathbf{y}}=0.05$. Further details regarding the forward measuring functions for each task and their respective hyperparameters are provided in the Appendix D.
Baselines. We compare our methods with the following baselines: Denoising Diffusion Restoration Models (DDRM) [2], Diffusion Posterior Sampling (DPS) [3], Denoising Diffusion Null-Space Model (DDNM) [31], and Plug-and-Play Alternating Direction Method of Multipliers (PnP-ADMM) [46] for pixel-based diffusion model experiments. We compare the latent diffusion version of our methods with Posterior Sampling with Latent Diffusion Models (PSLD) [27] and ReSample [25] that operates also in the latent space. Note that DDRM, DDNM, and PSLD are not able to handle nonlinear inverse problems. We also include Regularization by denoising diffusion process (RED-diff) [28] for nonlinear experiments.
[…]
Method SR (×4) Inpaint (Box) Inpaint (Random) Gaussian deblurring Motion deblurring
[…]
DAPS (ours) 0.177 29.07 0.133 24.07 0.098 31.12 0.165 29.19 0.157 29.66
DPS 0.260 24.38 0.198 23.32 0.193 28.39 0.211 25.52 0.270 23.14
DDRM 0.210 27.65 0.159 22.37 0.218 25.75 0.236 23.36 - -
DDNM 0.197 28.03 0.235 24.47 0.121 29.91 0.216 28.20 - -
PnP-ADMM 0.725 23.48 0.775 13.39 0.724 20.94 0.751 21.31 0.703 23.40
LatentDAPS (ours) 0.275 27.48 0.194 23.99 0.157 30.71 0.234 27.93 0.283 27.00
PSLD 0.287 24.35 0.158 24.22 0.221 30.31 0.316 23.27 0.336 22.31
ReSample 0.392 23.29 0.184 20.06 0.140 29.61 0.255 26.39 0.198 27.41
[…]
Method Phase retrieval Nonlinear deblurring High dynamic range
[…]
DAPS (ours) 0.121 (0.073) 30.72 (3.06) 0.155 (0.032) 28.29 (1.77) 0.162 (0.072) 27.12 (3.53)
DPS 0.410 (0.090) 17.64 (2.97) 0.278 (0.060) 23.39 (2.01) 0.264 (0.156) 22.73 (6.07)
RED-diff 0.596 (0.092) 15.60 (4.48) 0.160 (0.034) 30.86 (0.51) 0.258 (0.089) 22.16 (3.41)
LatentDAPS (ours) 0.199 (0.078) 29.16 (3.55) 0.235 (0.049) 28.11 (1.75) 0.223 (0.080) 25.94 (2.87)
ReSample 0.406 (0.224) 21.60 (8.10) 0.185 (0.039) 28.24 (1.69) 0.182 (0.085) 25.65 (3.57)
Table 3: Quantitative evaluation on ImageNet 256$\times$256. Performance comparison of different methods on various linear tasks in image domain.
Method SR (×4) Inpaint (Box) Inpaint (Random) Gaussian deblurring Motion deblurring
[…]
DAPS(ours) 0.276 25.89 0.214 21.43 0.135 28.44 0.253 26.15 0.196 27.86
DPS 0.354 23.92 0.309 19.78 0.326 24.43 0.360 21.86 0.357 21.46
DDRM 0.284 25.21 0.229 19.45 0.325 23.23 0.341 23.86 - -
DDNM 0.475 23.96 0.319 21.64 0.191 31.16 0.278 28.06 - -
PnP-ADMM 0.724 22.18 0.702 12.61 0.680 20.03 0.729 20.47 0.684 24.23
LatentDAPS(ours) 0.343 25.06 0.340 17.19 0.219 27.59 0.349 25.05 0.296 26.83
PSLD 0.360 25.42 0.465 20.10 0.337 31.30 0.390 25.86 0.511 20.85
ReSample 0.370 22.61 0.262 18.29 0.143 27.50 0.254 25.97 0.227 26.94
Table 4: Quantitative evaluation on ImageNet 256$\times$256. Performance comparison of different methods on various nonlinear tasks in the image domain. The mean and standard deviation are computed over $100$ images.
Method Phase retrieval Nonlinear deblurring High dynamic range
[…]
DAPS(ours) 0.254 (0.125) 25.78 (6.92) 0.169 (0.056) 27.73 (3.23) 0.175 (0.107) 26.30 (4.15)
DPS 0.447 (0.099) 16.81 (3.61) 0.306 (0.081) 22.49 (3.20) 0.503 (0.106) 19.23 (2.52)
RED-diff 0.536 (0.129) 14.98 (3.75) 0.211 (0.083) 30.07 (1.41) 0.274 (0.198) 22.03 (5.90)
LatentDAPS (ours) 0.361 (0.150) 20.54 (6.41) 0.314 (0.080) 25.34 (3.44) 0.269 (0.099) 23.64 (4.10)
ReSample 0.403 (0.174) 19.24 (4.21) 0.206 (0.057) 26.20 (3.71) 0.198 (0.089) 25.11 (4.21)
4.2 Main Results
We show quantitative results for the FFHQ dataset in Tables 1 and 2, and results for the ImageNet dataset in Tables 3 and 4. The qualitative comparisons are provided in Fig. 1. As shown in Table 1, our method achieves comparable or even better performance across all selected linear inverse problems. Moreover, our methods are remarkably stable when handling nonlinear inverse problems. For example, although existing methods such as DPS can recover high-quality images from phase retrieval measurements, they suffer from an extremely high failure rate, resulting in a relatively low PSNR and a high LPIPS. However, DAPS demonstrates superior stability across different samples and achieves a significantly higher success rate, resulting in a substantial improvement in PSNR and LPIPS as evidenced by the results in Table 2. Moreover, DAPS captures and recovers much more fine-grained details in measurements compared to existing baselines. We include more samples and comparisons in Appendix F for further illustration.
In addition, our method can generate diversified samples when the measurements contain less information. To illustrate this, we selected two tasks where the posterior distributions may have multiple modes: (1) super-resolution by a factor of 16, and (2) box inpainting with a box size of 192$\times$192. Some samples from DAPS are shown in Fig. 5, demonstrating DAPS’s ability to produce diverse samples while preserving the measurement information. We show more samples in Appendix F for further demonstration.
(a) Super resolution of a factor $16$
(b) Inpainting of 192$\times$192 box
Figure 6: Quantitative evaluations of image quality metrics and measurement error throughout sampling process. Here $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$ stands for the estimated mean of the Gaussian approximated distribution and ${\mathbf{x}}_{0\mid{\mathbf{y}}}\sim p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t},{\mathbf{y}})$ means the samples from measurement conditioned time-marginal. The measurement errors are normalized between $[0,1]$ for visualization.
4.3 Ablation Studies
Effectiveness of the number of function evaluations.
To better understand how the number of function evaluations (NFE) of the diffusion model influences the performance, we evaluate the performance of DAPS with different configurations. Recall that we use an ODE sampler in each inner loop to compute $\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})$, the total NFE for DAPS is the number of inner ODE steps times the number of noise annealing steps. We evaluate DAPS using NFE ranging from 50 to 4k, with configurations as specified in Section C.1. As indicated by Fig. 7, DAPS achieves relatively decent performance even with small NFE.
Table 5: Phase retrieval of different oversamples with DAPS.
Oversample 2.0 1.5 1.0 0.5 0.0
LPIPS 0.117 0.131 0.235 0.331 0.489
PSNR 30.26 29.17 24.87 21.60 16.02
More Discussion on Phase Retrieval.
Compared to the baselines, DAPS exhibits significantly better sample quality and stability in the phase retrieval task. Unlike other selected tasks, phase retrieval is more ill-posed, meaning that images with the same measurements can appear quite different perceptually. A demonstration of this is shown in Section F.1. To understand how this ill-posedness affects performance, we present results on $20$ images from FFHQ in Table 5 using different oversampling ratios in phase retrieval. These results further demonstrate the strength of DAPS in addressing complex, ill-posed inverse problems.
5 Conclusion
However, DAPS has its limitation and could be further improved in future work. For example, our method approximates $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})$ with a Gaussian distribution. This inaccurate approximation might lead to irrecoverable error in posterior sampling. Moreover, our method calls forward models multiple times while running Langevin dynamics, which increases the computation overhead when the forward model is complicated. Further discussions about limitations and future extensions are covered in Section C.2.
Acknowledgments
Appendix A Sampling with Latent Diffusion Models
@@
Remark. We can efficiently sample from $p({\mathbf{x}}_{0}\mid{\mathbf{z}}_{t_{1}},{\mathbf{y}})$ using similar strategies as in Section 3, i.e.,
@@
Both approaches are applicable for our posterior sampling algorithm. We summarize DAPS with latent diffusion models in Algorithm 2.
Algorithm 2 Decoupled Annealing Posterior Sampling with Latent Diffusion Models
@@
Appendix B Proof for Propositions

### v2

@@ A.1 Sampling with Hamiltonian Monte Carlo
         or HMC:
$(\hat{{\mathbf{x}}}_{0}^{(j+1)},{\mathbf{p}}^{(j+1)})\leftarrow\text{Hamiltonian-Dynamics}(\hat{{\mathbf{x}}}_{0}^{(j)},{\mathbf{p}}^{(j)}),$
         or Metropolis Hasting:
$\hat{{\mathbf{x}}}_{0}^{(j+1)}\leftarrow\text{Metropolis-Hasting}(\hat{{\mathbf{x}}}_{0}^{(j)})$
@@
A.2 Sampling with Metropolis Hasting algorithm
The Metropolis-Hastings algorithm [37] is a MCMC method used to sample from a target distribution $p({\mathbf{x}})$, especially when direct sampling is infeasible. It works by constructing a Markov chain whose stationary distribution corresponds to $p({\mathbf{x}})$. Here we discuss how to sample from $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t},{\mathbf{y}})$ using Metropolis Hasting under Gaussian approximation. We adopt the Gaussian kernel as the proposal distribution which is symmetry,
$q({\mathbf{x}}\to{\mathbf{x}}^{\prime})=\mathcal{N}({\mathbf{x}}^{\prime};{\mathbf{x}},\eta_{t}^{2}{\bm{I}})$ (12)
where $\eta_{t}$ is a hyperparameter to control the strength of the perturbation at diffusion time step $t$. Then we give the process of the Metropolis Hasting algorithm.
  1. 1.
    Start with the estimator from probability flow ODE $\hat{{\mathbf{x}}}_{0}^{(0)}=\hat{\mathbf{x}}_{0}({\mathbf{x}}_{x_{t}})$.
  2. 2.
    Perturb the current position by zero-centered Gaussian noise with standard deviation $\eta_{t}$, and number of steps $N$, for $j=1,\cdots,N$,
    ${\mathbf{x}}_{prop}^{(j)}=\hat{\mathbf{x}}_{0}^{(j)}+\eta_{t}\epsilon,\quad\epsilon\sim\mathcal{N}({\bm{0}},{\bm{I}})$ (13)
  3. 3.
    Update position by,
    $\hat{\mathbf{x}}_{0}^{(j+1)}=\text{Metropolis-Hasting}(\hat{{\mathbf{x}}}_{0}^{(j)})=\begin{cases}{\mathbf{x}}_{prop}^{(j)}\quad&\text{w.p. }\,\alpha(\hat{\mathbf{x}}_{0}^{(j)}\to{\mathbf{x}}_{prop}^{(j)})\\ \hat{\mathbf{x}}_{0}^{(j)}\quad&\text{otherwise}\end{cases},$ (14)
    where the acceptance probability $\alpha(\hat{\mathbf{x}}_{0}^{(j)}\to{\mathbf{x}}_{prop}^{(j)})$ is given by,
@@ A.2 Sampling with Metropolis Hasting algorithm
    $\alpha(\hat{\mathbf{x}}_{0}^{(j)}\to{\mathbf{x}}_{prop}^{(j)})=\min\left(1,\dfrac{p({\mathbf{x}}_{prop}^{(j)}\mid{\mathbf{x}}_{t},{\mathbf{y}})}{p(\hat{\mathbf{x}}_{0}^{(j)}\mid{\mathbf{x}}_{t},{\mathbf{y}})}\right)\approx\min\left(1,\dfrac{\mathcal{N}({\mathbf{x}}_{prop}^{(j)};{\mathbf{x}}_{t},\sigma_{t}^{2}{\bm{I}})\cdot\mathcal{N}({\mathbf{y}};\mathcal{A}({\mathbf{x}}_{prop}^{(j)}),\beta_{\mathbf{y}}^{2}{\bm{I}})}{\mathcal{N}(\hat{\mathbf{x}}_{0}^{(j)};{\mathbf{x}}_{t},\sigma_{t}^{2}{\bm{I}})\cdot\mathcal{N}({\mathbf{y}};\mathcal{A}(\hat{\mathbf{x}}_{0}^{(j)}),\beta_{\mathbf{y}}^{2}{\bm{I}})}\right)$ (15)
Empirically, Metropolis Hasting usually performs worse than Langevin dynamics and HMC and less efficient but more flexible to tasks that don’t have access to the gradient of the data likelihood, $\nabla_{{\mathbf{x}}_{0}}\log p({\mathbf{y}}\mid{\mathbf{x}}_{0})$, i.e. the forward function $\mathcal{A}$ is non-differentiable or its gradient is inaccessible. We will discuss more about such tasks in Appendix C.
Appendix B DAPS with Latent Diffusion Models
@@
Remark. We can efficiently sample from $p({\mathbf{x}}_{0}\mid{\mathbf{z}}_{t_{1}},{\mathbf{y}})$ using similar strategies as in Sec. 3, i.e.,
@@
Both approaches are applicable to our posterior sampling algorithm. We summarize DAPS with latent diffusion models in Algorithm 2. It is worth mentioning that when employing the Gaussian approximation for $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})$, pixel-space Langevin dynamics typically results in higher approximation errors compared to latent-space Langevin dynamics but offers significantly faster computation. A practical approach is to utilize pixel-space Langevin dynamics during the early stages (i.e., for $N_{A}\geq i>M$) to balance approximation accuracy with efficiency, transitioning to latent-space Langevin dynamics in the later stages (i.e., for $M\geq i\geq 1$) to achieve improved sample quality, where $M$ is a hyperparameter that trade-off efficiency and approximation accuracy. For simplicity, we set $M=N_{A}$ throughout our experiments, and leave the exploration of this hyperparameter as a future direction.
Algorithm 2 Latent Diffusion Decoupled Annealing Posterior Sampling (LatentDAPS) with Langevin Dynamics
@@
B.1 Sampling with Large-scaled Text-conditioned Latent Diffusion Models
Large-scale text-conditioned latent diffusion models (LDMs) [39, 40] provide a diverse and powerful prior for solving inverse problems. Given a text prompt ${\mathbf{c}}$, our goal is to sample from $p({\mathbf{x}}_{0}\mid{\mathbf{c}},{\mathbf{y}})\propto p({\mathbf{y}}\mid{\mathbf{x}}_{0})p({\mathbf{x}}_{0}\mid{\mathbf{c}})$, where $p({\mathbf{x}}_{0}\mid{\mathbf{c}})$ is modeled by the large-scale text-conditioned LDMs. To achieve this, we can utilize Algorithm 2 for posterior sampling.
To evaluate the performance of DAPS with text-conditioned LDMs, we use Stable Diffusion v1.5 [40] and assess it on the same eight tasks from the FFHQ 256 dataset, as described in [42, 43]. The sampling process is enhanced using classifier-free guidance for text conditioning with a guidance scale of $7.5$. The quantitative results are shown in Tab. 4, and the qualitative results are presented in Fig. 8. A significant advantage of text-conditioned LDMs is their ability to flexibly control posterior sampling using text prompts. As illustrated in Fig. 9, text prompts enable effective exploration of different modes in the posterior distribution. Finally, we summarize the text prompts used for each task in Tab. 5.
[…]
Method Super Resolution 4× Inpaint (Box) Inpaint (Random) Gaussian deblurring Motion deblurring Phase retrieval Nonlinear deblurring High dynamic range
[…]
LatentDAPS (SD-v1.5, ours) 0.109 31.43 0.133 $23.33$ 0.064 34.60 0.160 30.73 $0.101$ $33.84$ $0.256$ $27.47$ $0.116$ $31.67$ $0.186$ $24.52$
PSLD (SD-v1.5) $0.201$ $30.73$ $0.167$ $0.096$ $30.31$ $0.221$ $30.10$
[…]
Figure 9: Exploring different modes in the posterior distribution by text prompts. Multiple sampling results using LatentDAPS (SD v1.5) were obtained for the super resolution 8× problem on a 512$\times$512 natural image. Text prompts provided additional control over the sampled results..
Table 5: The text prompts used in LatentDAPS (SD v1.5).
Text Prompts
FFHQ Evaluation A natural looking human face.
Teaser Fig. 1 Blue butterfly on white flower, green blurred background.
Juicy burger with toppings, fresh fries, blurred restaurant background.
Sunlit mountain reflected in a serene lake, surrounded by trees.
Appendix C DAPS with Discrete Diffusion Models
Discrete diffusion models [3, 6, 33] are designed to generate categorical data over a finite support. Similar to diffusion models in the continuous space, discrete diffusion models evolve a family of distributions $p_{0}({\mathbf{x}}),\dots,p_{T}({\mathbf{x}})$ according to a continuous-time Markov chain. Specifically, the forward diffusion process is defined as
$\frac{\mathrm{d}p_{t}}{\mathrm{d}t}={\bm{Q}}_{t}p_{t},$ (24)
where ${\bm{Q}}_{t}$ are predefined transition matrices, such that $p_{t}$ converges to a simple distribution like uniform distribution or a special “mask” state, as $t\to\infty$. To reverse the forward process, it suffices to learn the “concrete score”, $s({\mathbf{x}},t)=\Big{[}\frac{p_{y}({\mathbf{y}})}{p_{t}({\mathbf{x}})}\Big{]}_{{\mathbf{y}}\neq{\mathbf{x}}}$. The reverse diffusion process is given by
$\frac{\mathrm{d}p_{T-t}}{\mathrm{d}t}=\overline{\bm{Q}}_{T-t}p_{T-t},$ (25)
where $\overline{\bm{Q}}_{t}({\mathbf{y}},{\mathbf{x}})=s({\mathbf{x}},t){\bm{Q}}_{t}({\mathbf{x}},{\mathbf{y}})$ and $\overline{\bm{Q}}_{t}({\mathbf{x}},{\mathbf{x}})=-\sum_{y\neq x}\overline{\bm{Q}}_{t}({\mathbf{y}},{\mathbf{x}})$.
Given the similarity of continuous and discrete diffusion models, we find that DAPS can be extended to perform posterior sampling with discrete diffusion models. Instead of making the Gaussian approximation as in continuous diffusion models, we approximate $p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})$ with an exponential distribution over Hamming distance, i.e.,
$p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t})\approx\exp(-\|{\mathbf{x}}_{0}-\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t})\|_{0}/r_{t}),$ (26)
where $r_{t}$ is determined heuristically. We put the DAPS sampling algorithm for discrete diffusion models as in Algorithm 3.
Algorithm 3 Decoupled Annealing Posterior Sampling (DAPS) with Discrete Diffusion Models
Score model ${\bm{s}}_{{\bm{\theta}}}$, measurement ${\mathbf{y}}$, $(t_{i})_{i\in\{0,\dots,N_{A}\}}$.
Sample ${\mathbf{x}}_{T}\sim p_{T}$.
for $i=N_{A},N_{A}-1,\dots,1$ do
     Compute $\hat{\mathbf{x}}_{0}^{(0)}=\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t_{i}})$ by solving the reverse continuous-time Markov chain in Eq. 25 with ${\bm{s}}_{\bm{\theta}}$
     for $j=0,\dots,N-1$ do
         Metropolis Hasting
$\hat{\mathbf{x}}_{0}^{(j+1)}\sim p({\mathbf{x}}_{0}\mid{\mathbf{x}}_{t},{\mathbf{y}})\approx p({\mathbf{y}}\mid{\mathbf{x}}_{0})\exp(\|{\mathbf{x}}_{0}-\hat{\mathbf{x}}_{0}({\mathbf{x}}_{t_{i}})\|_{0}/r_{t_{i}}).$
     end for
     Sample ${\mathbf{x}}_{t_{i-1}}\sim p({\mathbf{x}}_{t_{i-1}}\mid\hat{\mathbf{x}}_{0}^{(N)})$ following Eq. 24.
end for
Return ${\mathbf{x}}_{0}$
Experiments on discretized MNIST dataset.
We conduct experiments on the discretized MNIST dataset to demonstrate how DAPS can be applied to inverse problems on categorical data. We first discretize and flatten MNIST data to binary strings. We consider two inverse problems: 1) inpainting and 2) XOR operator. We mask out $50\%$ pixels for the inpainting task. For the XOR task, we draw $50\%$ random pairs from the MNIST binary strings and compute XOR over all the pairs, which serves as a highly nonlinear test case. We compare DAPS with best-of-N samples, and a recent work on discrete diffusion posterior sampling (SVDD-PM).
As shown in Tab. 6, DAPS with discrete diffusion models is able to achieve very high classification accuracy on both inverse problem tasks, outperforming both baselines by a large margin. This suggests the effectiveness of DAPS in solving inverse problems for categorical data. We leave further exploration of DAPS in discrete-state space for future work.
Table 6: Quantitative results on discretized MNIST on two discrete inverse problems. Both SVDD-PM and best-of-N use 20 particles for sampling. The classification accuracy is computed using a simple ConvNet model trained using MNIST training dataset.
Inpainting XOR
[…]
DAPS $\textbf{18.82}_{\pm 2.03}$ 97.0 $\textbf{20.50}_{\pm 6.40}$ 98.0
SVDD-PM $11.84_{\pm 2.56}$ $38.0$ $13.00_{\pm 2.88}$ $59.0$
Best-of-N $10.56_{\pm 1.11}$ $37.0$ $10.50_{\pm 1.13}$ $36.0$
Appendix D Proof for Propositions